A 15.6 cm -diameter coil consists of 27 turns of circular copper wire 2.6 mm in diameter. A uniform magnetic field, perpendicular to the plane of the coil, changes at a rate of 7.90×10−3 T/s . (PLEASE show step by step and don't write down random numbers without showing how you got it. Thank you!)

A. Determine the current in the loop.

B. Determine the rate at which thermal energy is produced.
Answer #1

area of loop

a = pi/4 d^2 = 3.14 / 4* (0.156^2) = 0.0191 m^2

total lemgth of wire

L = 27* pi d = 13.226 m

area of cross section of wire

A = 3.14/ 4* (2.6*10^-3)^2 = 5.307*10^-6 m^2

Resistance of wire

R = rho * L/ A

R = 1.72*10^-8* 13.226 / (5.307*10^-6)

R = 0.0429 ohm

now

emf induced,

e = N a dB/dt

i = e / R = 27 * 0.0191* 7.9*10^-3 / 0.0429

i = 0.095 A

=========

rate of thermal energy

E = i^2 R = 0.095^2* 0.0429= 3.8745*10^-4 W

========

Comment in case any doubt, will reply for sure.. Goodluck